OpenStudy (anonymous):

Find the values of x in the following equation: |-5x + -2| = 12

OpenStudy (anonymous):

| something| = 12 Two numbers have an absolute value of 12 (-12 and 12) we set the inside equal to both of those numbers -5x-2 = -12 and -5x-2 = 12 Solve for x. I'm getting x = 2 and x = -14/5 Hope that helps.

OpenStudy (anonymous):

|-5x+-2| = 12 |-(5x+2)|=12 those absolute they be gone 5x+2=12 then x = 2